# Language Model Evolutionary Algorithms for Recommender Systems: Benchmarks and Algorithm Comparisons

### v1

@@ Abstract
In the evolutionary computing community, the remarkable language-handling capabilities and reasoning power of large language models (LLMs) have significantly enhanced the functionality of evolutionary algorithms (EAs), enabling them to tackle optimization problems involving structured language or program code. Although this field is still in its early stages, its impressive potential has led to the development of various LLM-based EAs. To effectively evaluate the performance and practical applicability of these LLM-based EAs, benchmarks with real-world relevance are essential. In this paper, we focus on LLM-based recommender systems (RSs) and introduce a benchmark problem set, named RSBench, specifically designed to assess the performance of LLM-based EAs in recommendation prompt optimization. RSBench emphasizes session-based recommendations, aiming to discover a set of Pareto optimal prompts that guide the recommendation process, providing accurate, diverse, and fair recommendations. We develop three LLM-based EAs based on established EA frameworks and experimentally evaluate their performance using RSBench. Our study offers valuable insights into the application of EAs in LLM-based RSs. Additionally, we explore key components that may influence the overall performance of the RS, providing meaningful guidance for future research on the development of LLM-based EAs in RSs.
@@ I Introduction
Evolutionary algorithms (EAs), inspired by natural selection and genetic principles, constitute a powerful class of optimization techniques [1]. Unlike traditional methods, EAs do not rely on gradient information, making them highly suitable for tackling complex black-box optimization problems that commonly arise in scientific [2] and engineering [3] fields. Moreover, EAs are particularly effective for multiobjective optimization problems, where multiple conflicting objectives require simultaneous optimization [4]. Their population-based search strategy, which leverages implicit parallelism, allows EAs to explore diverse solution spaces within a single optimization run. This approach yields a set of optimized solutions—referred to as nondominated or Pareto optimal solutions—that align with the Pareto front (PF), representing the best achievable performance trade-offs in the objective space. Meanwhile, diversity-preserving mechanisms within EAs further enable the discovery of solutions that are well-distributed along the PF, allowing decision-makers to visualize the trade-offs and select the most preferred solution(s) a posteriori [5].
@@
Large language models (LLMs)[6], such as ChatGPT[7], have significantly advanced language processing technology, exhibiting exceptional performance in natural language processing tasks and fostering transformative applications across fields such as mechanical engineering [8], management [9], and chemistry [10]. This advancement has created promising opportunities to enhance EAs [11]. On one hand, the black-box optimization capabilities of EAs are well-suited for LLM-based applications. Due to the high computational demands of LLM training and inference, deployment is largely restricted to corporations like OpenAI, Google, and Baidu, which often limit access to model parameters for security and privacy, providing only API-based querying [11]. Consequently, users often treat LLMs as black-box systems, and EAs, as classic and effective black-box optimization algorithms, present a promising approach for refining LLM applications [12]. On the other hand, the generative capabilities of LLMs can help expand the functional scope of traditional EAs, which typically represent solutions as low-dimensional continuous or discrete vectors [13]. Integrating LLMs enables EAs to process language and code, allowing them to tackle complex tasks such as prompt optimization [14], neural architecture search [15], and code generation [16]. Consequently, recent research has increasingly explored hybrid approaches that combine EAs with LLMs, applying these techniques to both traditional optimization challenges and novel application domains.
@@
Recommender systems (RSs) play an increasingly crucial role on modern online platforms [17], and the emergence of LLMs also has introduced new opportunities for their development and enhancement [18]. In the early stages, language models such as BERT [19] were used to embed language elements from the recommendation process into suitable vectors, supporting the training of more accurate recommendation models [20]. As LLMs evolved, pre-training [21] and fine-tuning [22] paradigms were developed, leading to the creation of even more powerful recommendation performance [23]. Recently, with the advent of models like ChatGPT [7], the advanced language processing capabilities of LLMs enable the creation of robust RSs through simple prompts [24, 25, 26] or in-context learning [27, 28].
@@
In prompt-based LLM-driven RSs, the design of prompts plays a pivotal role in shaping system performance [26]. As a result, finding the optimal prompt becomes a critical priority when developing such systems. Additionally, in RSs, it is essential for recommendations not only to align with user-specific interests but also to consider factors such as diversity and fairness to enhance overall user satisfaction [29, 30]. Multiobjective optimization serves as an ideal theoretical framework for this purpose [31]. By leveraging this approach, the system can provide users with a set of Pareto-optimal solutions that balance criteria like recommendation accuracy, diversity, and fairness, thus allowing users to select options that best meet their personalized preferences. Whether addressing black-box prompt optimization or multiobjective optimization, EAs have demonstrated remarkable capabilities. Consequently, EAs hold significant potential for application in LLM-based RSs. However, to the best of our knowledge, there has no research, both in terms of benchmark problems and algorithms, on applying EAs to LLM-based RSs. To bridge this gap, this paper explores the application of language model based EAs in RSs. The key contributions of this paper are mainly focus on the follow aspects:
  • We introduce a benchmark problem set, named RSBench, specifically designed for EAs on LLM-based RSs. RSBench focuses primarily on session-based recommendation [26], which aims to predict next promising items for interaction based on short, anonymous behavior sessions. The goal is to discover a set of Pareto optimal prompts that can guide the recommendation process, allowing users to freely choose recommendations based on their preferences regarding accuracy, diversity, and fairness, balancing two or three of these aspects.
@@
  • A set of LLM-based multiobjective EAs are developed to solve the RSBench problems. Following the frameworks of three representative multiobjective EAs—NSGA-II, MOEA/D, and IBEA—LLM-based operations are integrated into these classical algorithms to support their application in solving RSBench problems. Their performance is then evaluated experimentally, providing valuable insights and guidance for the application of EAs in LLM-based RSs.
@@
  • We further investigate key components that may impact the performance of the final RS, including the effectiveness of the LLM-based initialization and the LLM-based offspring generation with the supporting of the crossover and mutation operators. These investigations are expected to provide valuable insights and guidance for future research on LLM-based EAs in RSs.
  • Finally, with regard to LLM-based EAs for RSs, we discuss promising research topics and pressing issues that warrant future exploration, providing guidance for subsequent research.
@@
The remainder of this paper is organized as follows: Section II provides a review of related work and background concepts. Section III introduces RSBench. In Section IV, we present the three LLM-based multiobjective EAs, based on the classical NSGA-II, MOEA/D, and IBEA frameworks. Section V discusses the systematic experimental investigations. In Section VI, we explore promising research directions based on RSBench and highlight pressing issues for future exploration. Finally, Section VII concludes the paper.
@@ III-B2 Objective Functions
    where $T^{k}_{llm}(\textbf{X}_{s}^{(l)},\mathcal{Y}_{c}^{(l)})$ denotes the item ranked $k$ by the LLM-based RS, $C[\cdot]$ represents the category set associated with an item, and ||$\cdot$|| ⋅ | denotes the size of a set. The formula first calculates the ratio of the number of unique categories among the top-$K$ items to the sum of the categories of each top-$K$ item for each recommendation sample. It then computes the average ratio across all N recommendation samples. This formula reflects diversity by assigning smaller ratios to cases with high category overlap among top-$K$ items, which implies reduced diversity. Hence, higher values of $F_{div}$ indicate greater diversity in the recommendations.
@@ III-B3 Instances of RSBench
Based on the definitions of the decision variables and objective functions provided above, we introduce nine instances of the RSBench problem set. To obtain the objective function values, datasets containing customer consumption and click data are required. In this benchmark, we use three real-world datasets from different domains. Specifically, MovieLens-1M (ML-1M) [^1] https://grouplens.org/datasets/movielens/ is a movie dataset containing customer ratings for various films. The “Games” dataset is a subcategory from the Amazon dataset [64], which includes customer ratings for video games. The “Bundle” dataset [69] consists of session data from three Amazon subcategories: Electronics, Clothing, and Food. The statistical features of these datasets are summarized in Table I. For the ML-1M and Games datasets, we chronologically order the rated items for each user and divide them into sessions by day. All datasets are split into training and validation sets with a ratio of 8:2.
@@ IV-A1 Initialization of the Recommendation Prompt Population
While the robust capabilities of LLMs enable them to excel at various downstream tasks, their performance is highly dependent on the quality of the prompts provided. Over the past year, numerous prompt structures, such as chain-of-thought [70] and tree-of-thought [71], have been proposed to improve LLM performance. In this paper, to expedite the discovery of high-quality prompts, we introduce a well-designed recommendation prompt as the example based on the chain-of-thought approach. The details of this example recommendation prompt are presented in Fig. 3. The initial population of recommendation prompts is subsequently generated using the LLM based on the example recommoned prompt, with the initialization prompt shown in Fig. 4. In this manner, $N$ recommendation prompts are produced to form the initial population.
@@ IV-A2 Offspring Recommendation Prompts Generation
In traditional EAs, crossover and mutation operators are used to generate offspring solutions. However, these operators are typically designed for continuous or discrete vectors and are not well-suited for handling natural language, such as recommendation prompts. Recent studies have demonstrated that, with appropriate prompts, LLMs can themselves serve as effective crossover and mutation operators for natural language processing tasks [12]. Following this approach, we employ an LLM-based crossover and mutation mechanism in this work. The prompt used for the LLM-based crossover and mutation operator is illustrated in Fig. 3. By providing two parent recommendation prompts to the LLM-based operator, an offspring recommendation prompt is generated. This LLM-based operator thus facilitates the generation of an offspring population from the parent population.
@@ IV-B1 LLM-NSGA-II
0:  The nondominated recommendation prompt set $\mathcal{P}$;
1:  Generate the initial population $\mathcal{P}=\{\textbf{p}_{1},\ldots,\textbf{p}_{N}\}$ by using the initial operations introduced in Section IV-A1;
2:  Evaluate the initial population by using the objective functions;
3:  $t\leftarrow 0$
4:  while $t<t_{max}$ do
5:     Generate an offspring population of recommendation prompts $\mathcal{Q}=\{\textbf{q}_{1},\ldots,\textbf{q}_{N}\}$ according to the operations introduced in Section IV-A2;
6:     Evaluate the offspring population by using the objective functions;
7:     $\mathcal{P}\leftarrow\mathcal{P}\cup\mathcal{Q}$;
8:     Sorting the recommendation prompts in $\mathcal{P}$ according to the nondominated sorting and the crowd distance, and retain only the top $N$ recommendation prompts;
9:     $t\leftarrow t+1$;
10:  end while
11:  return  A population of nondominated recommendation prompts $\mathcal{P}$.
@@ IV-B2 LLM-MOEA/D
0:  The nondominated recommendation prompt set $\mathcal{P}$;
1:  Generate the neighborhood set $\mathcal{B}_{i}$ for the weight vector $\textbf{w}_{i}$, where $i\in\{1,\ldots,N\}$;
2:  Generate the initial population $\mathcal{P}=\{\textbf{p}_{1},\ldots,\textbf{p}_{N}\}$ by using the initial operations introduced in Section IV-A1;
3:  Evaluate the initial population by using the objective functions;
4:  Initial the external population $\mathcal{E}$ and the reference point $\textbf{z}=(z_{1},...,z_{m})^{T}$;
5:  $t\leftarrow 0$
6:  while $t<t_{max}$ do
7:     for $i=\{1,\ldots,N\}$ do
8:        Randomly select two indexes $k$ and $l$ from $\mathcal{B}_{i}$, and generate a new prompt $\textbf{p}^{new}$ according to $\textbf{p}^{k}$ and $\textbf{p}^{l}$ by using the LLM-based offspring generation introduced in Section IV-A2;
9:        Evaluate $\textbf{p}^{new}$ by using the objective functions;
10:        Update the reference point z, the population $\mathcal{P}$ and the external population $\mathcal{E}$ following the MOEA/D framework;
11:     end for
12:     $t\leftarrow t+1$;
13:  end while
14:  return  An external population of nondominated recommendation prompts $\mathcal{E}$.
Decomposition-based evolutionary multiobjective optimization is a classical paradigm in multiobjective optimization. LLM-MOEA/D follows the decomposition-based method, with the pseudo-code presented in Algorithm 2. As in the MOEA/D framework, a neighborhood set $\mathcal{B}^{i}$ is initialized for each weight vector $\textbf{w}^{i}$. The initial population $\mathcal{P}$ is generated using the LLM-based initialization described in Section IV-A1, and all recommendation prompts in $\mathcal{P}$ are evaluated using the objective functions. Based on these evaluations, the external population $\mathcal{E}$ and the reference point z are initialized. During the iteration process, for each weight vector, an offspring recommendation prompt $\textbf{p}^{new}$ is produced using the LLM-based offspring generation method detailed in Section IV-A2. After evaluating $\textbf{p}^{new}$, the reference point z, the population $\mathcal{P}$, and the external population $\mathcal{E}$ are updated in accordance with the MOEA/D framework [36]. Once the termination conditions are satisfied, the external population $\mathcal{E}$ of nondominated recommendation prompts is returned.

### v3

@@ Abstract
In the evolutionary computing community, the remarkable language-handling capabilities and reasoning power of large language models (LLMs) have significantly enhanced the functionality of evolutionary algorithms (EAs), enabling them to tackle optimization problems involving structured language or program code. Although this field is still in its early stages, its impressive potential has led to the development of various LLM-based EAs. To effectively evaluate the performance and practical applicability of these LLM-based EAs, benchmarks with real-world relevance are essential. In this paper, we focus on LLM-based recommender systems (RSs) and introduce a benchmark problem set, named RSBench, specifically designed to assess the performance of LLM-based EAs in recommendation prompt optimization. RSBench emphasizes session-based recommendations, aiming to discover a set of Pareto optimal prompts that guide the recommendation process, providing accurate, diverse, and fair recommendations. We develop three LLM-based EAs based on established EA frameworks and experimentally evaluate their performance using RSBench. Our study offers valuable insights into the application of EAs in LLM-based RSs. Additionally, we explore key components that may influence the overall performance of the RS, providing meaningful guidance for future research on the development of LLM-based EAs in RSs. The source code of the proposed RSBench can be found at https://github.com/LiuJ-2023/RSBench/tree/main.
@@ I Introduction
Evolutionary algorithms (EAs), inspired by natural selection and genetic principles, constitute a powerful class of optimization techniques [1]. Unlike traditional methods, EAs do not rely on gradient information, making them highly suitable for tackling complex black-box optimization problems that commonly arise in scientific [2] and engineering [3] fields. Moreover, EAs are particularly effective for multiobjective optimization problems, where multiple conflicting objectives require simultaneous optimization [4]. Their population-based search strategy, which leverages implicit parallelism, allows EAs to explore diverse solution spaces within a single optimization run. This approach yields a set of optimized solutions—referred to as nondominated or Pareto optimal solutions—that align with the Pareto front (PF), representing the best achievable performance trade-offs in the objective space. Meanwhile, diversity-keeping mechanisms within EAs facilitate the identification of solutions that are well-distributed along the PF, resulting in a more accurate approximation of the PF. [5].
@@
Large language models (LLMs)[6, 7] have significantly advanced language processing technology, exhibiting exceptional performance in natural language processing tasks and fostering transformative applications across fields such as mechanical engineering [8], management [9], and chemistry [10]. This advancement has created promising opportunities to enhance EAs [11]. On one hand, the black-box optimization capabilities of EAs are well-suited for LLM-based applications. Due to the high computational demands of LLM training and inference, deployment is largely restricted to corporations like OpenAI, Google, and Baidu, which often limit access to model parameters for security and privacy, providing only API-based querying [11]. Consequently, users often treat LLMs as black-box systems, and EAs, as classic and effective black-box optimization algorithms, present a promising approach for refining LLM applications [12]. On the other hand, the generative capabilities of LLMs can help expand the functional scope of traditional EAs, which typically represent solutions as low-dimensional continuous or discrete vectors [13]. Integrating LLMs enables EAs to process language and code, allowing them to tackle complex tasks such as prompt optimization [14], neural architecture search [15], and code generation [16]. Consequently, recent research has increasingly explored hybrid approaches that combine EAs with LLMs, applying these techniques to both traditional optimization challenges and novel application domains.
@@
Recommender systems (RSs) play an increasingly crucial role on modern online platforms [17], and the emergence of LLMs also has introduced new opportunities for their development and enhancement [18]. Traditional RS algorithms, such as collaborative filtering [19], primarily rely on ID-based information while overlooking the rich textual data generated through user interactions on the platform [20]. In fact, leveraging these textual data can help mitigate inherent limitations in traditional recommendation algorithms, such as the cold-start problem in collaborative filtering [21], while also capturing users’ current preferences and intentions [22]. The powerful text-processing capabilities of LLMs further enable RSs to effectively utilize textual data, leading to more sophisticated and context-aware recommendations. In the early stages, language models such as BERT [23] were used to embed language elements from the recommendation process into suitable vectors, supporting the training of more accurate recommendation models [24]. As LLMs evolved, pre-training [25] and fine-tuning [26] paradigms were developed, leading to the creation of even more powerful recommendation performance [27]. Recently, with the advent of models like ChatGPT [7], the advanced language processing capabilities of LLMs enable the creation of robust RSs through simple prompts [28, 29, 22] or in-context learning [30, 31].
@@
In prompt-based LLM-driven RSs, the design of prompts plays a pivotal role in shaping system performance [22]. As a result, finding the optimal prompt becomes a critical priority when developing such systems. Additionally, in RSs, it is essential for recommendations not only to align with user-specific interests but also to consider factors such as diversity and fairness to enhance overall user satisfaction [32, 33]. Multiobjective optimization serves as an ideal theoretical framework for this purpose [34]. By leveraging this approach, the system can provide users with a set of Pareto-optimal solutions that balance criteria like recommendation accuracy, diversity, and fairness, thus allowing users to select options that best meet their personalized preferences. Whether addressing black-box prompt optimization or multiobjective optimization, EAs have demonstrated remarkable capabilities. Consequently, EAs hold significant potential for application in LLM-based RSs. However, to date, there are almost no benchmark problems specifically addressing the application of LLM-based EAs to LLM-based RSs. To bridge this gap, this paper explores the application of language model based EAs in RSs. The key contributions of this paper are mainly focus on the follow aspects:
  • In this work, we introduce RSBench, a benchmark specifically designed for LLM-based RSs. RSBench provides a convenient and accessible testbed that enables EA researchers to explore, validate, and advance evolutionary algorithms in the context of LLM-based recommendation. By bridging this gap, RSBench encourages stronger engagement between the EA and RS communities and promotes deeper investigation into the potential and practical value of EA for real-world RS challenges. To support reproducibility and further research, we release the accompanying code at https://github.com/LiuJ-2023/RSBench, serving as a foundational resource for understanding and advancing this field.
@@
  • LLM-based initialization and offspring generation are designed and incorporated into three representative multiobjective EAs—  i.e., NSGA-II, MOEA/D, and IBEA—to support their application in solving RSBench problems. Their performance is then evaluated experimentally, providing valuable insights and guidance for the application of LLM-based EAs in LLM-based RSs.
@@
  • We further investigate key components that may impact the performance of the final RS, including the effectiveness of the LLM-based initialization and the LLM-based offspring generation with the supporting of the crossover and mutation operators. These investigations are expected to point out challenges and potential research questions for future research on LLM-based EAs in RSs.
  • RSBench can also serve as a test function suite for prompt optimization in application domains. It incorporates characteristics such as expansionality, uncertainty, and language-space search, thereby reflecting many of the distinctive properties of prompt optimization problems at the LLM application level.
@@
The remainder of this paper is organized as follows: Section II provides a review of related work and background concepts. Section III introduces RSBench. In Section IV, we present the three LLM-based multiobjective EAs, based on the classical NSGA-II, MOEA/D, and IBEA frameworks. Section V discusses the systematic experimental investigations, which aims to provide insights and guidance for researchers in the related area. In Section VI, we explore promising research directions based on RSBench and highlight pressing issues for future exploration. Finally, Section VII concludes the paper.
@@ III-B2 Objective Functions
    where $T^{k}_{llm}(\textbf{X}_{s}^{(l)},\mathcal{Y}_{c}^{(l)})$ denotes the item ranked $k$ by the LLM-based RS, $C[\cdot]$ represents the category set associated with an item, and $|\cdot|$ denotes the size of a set. The formula first calculates the ratio of the number of unique categories among the top-$K$ items to the sum of the categories of each top-$K$ item for each recommendation sample. It then computes the average ratio across all N recommendation samples. This formula reflects diversity by assigning smaller ratios to cases with high category overlap among top-$K$ items, which implies reduced diversity. Hence, higher values of $F_{div}$ indicate greater diversity in the recommendations.
@@ III-B3 Instances of RSBench
Based on the definitions of the decision variables and objective functions provided above, we introduce nine instances of the RSBench problem set. To obtain the objective function values, datasets containing customer consumption and click data are required. In this benchmark, we use three real-world datasets from different domains. Specifically, MovieLens-1M (ML-1M) [^1] https://grouplens.org/datasets/movielens/ is a movie dataset containing customer ratings for various films. The “Games” dataset is a subcategory from the Amazon dataset [67], which includes customer ratings for video games. The “Bundle” dataset [72] consists of session data from three Amazon subcategories: Electronics, Clothing, and Food. The statistical features of these datasets are summarized in Table I. For the ML-1M and Games datasets, we chronologically order the rated items for each user and divide them into sessions by day.
@@
III-B4 Challenges of the RSBench Problems
In general, we consider that the RSBench problems mainly contains the following three challenges:
  • High Computational Costs: As shown in (2), (3), and (4), all objective values are evaluated using the Monte Carlo method. Consequently, computing the objective function requires testing each sample individually and repeatly. This frequent interaction with APIs incurs significant computational costs in terms of time and token consumption, making RSBench problems inherently expensive. When applying EAs to RSBench, this computational expense must be carefully considered to ensure efficient optimization.
  • Uncertainty: Additionally, the evaluation of objective functions introduces uncertainty in assessing the true performance of recommendation prompts. This uncertainty arises from three primary sources. First, the Monte Carlo evaluation may deviate from the actual performance of a prompt, introducing potential bias. Second, the stochastic nature of LLM inference leads to variability across repeated trials. Third, users themselves may not always exhibit clear or consistent preferences, further embedding uncertainty into RSBench. Such uncertainties can misguide the evolutionary process, ultimately hindering convergence and degrading overall performance.
  • Searching in the Language Space: In the RSBench problem, our objective is to identify an optimal recommendation prompt to guide the LLM in completing the recommendation task. Unlike conventional EAs, which search for optimal solutions in a fixed-dimensional vector space, RSBench requires searching for optimal solutions in a language space. This unique search space introduces two key challenges. First, since the solutions are expressed in natural language, they must maintain linguistic coherence, which imposes implicit constraints on word selection and arrangement. Second, as natural language does not have a fixed structure or length, the optimization process must operate within a variable-length search space, requiring adaptive strategies to effectively explore and refine solutions.
@@ IV-A1 Initialization of the Recommendation Prompt Population
While the robust capabilities of LLMs enable them to excel at various downstream tasks, their performance is highly dependent on the quality of the prompts provided. Over the past year, numerous prompt structures, such as chain-of-thought [73] and tree-of-thought [74], have been proposed to improve LLM performance. In this paper, to expedite the discovery of high-quality prompts, we introduce a well-designed recommendation prompt as the example based on the chain-of-thought approach. The details of this example recommendation prompt are presented in Fig. 3. This example recommendation prompt guides the recommendation process through four structured steps. First, it identifies meaningful item combinations within the session to capture contextual relationships. Second, it infers the user’s interactive intent based on these combinations. Third, it selects the most relevant intent that best represents the user’s preferences. Finally, it re-ranks the candidate items accordingly. By systematically breaking down user interactions, this prompt ensures that recommendations are context-aware and preference-driven. The initial population of recommendation prompts is subsequently generated using the LLM based on the example recommoned prompt, with the initialization prompt shown in Fig. 4. In this manner, $N$ recommendation prompts are produced to form the initial population. Such an initialization process ensures, on the one hand, that the recommendation prompts in the generated initial population are logically structured to guide the LLM in effectively completing the recommendation task. On the other hand, it also introduces a certain degree of diversity within the population.
@@ IV-A2 Offspring Recommendation Prompts Generation
In traditional EAs, crossover and mutation operators are used to generate offspring solutions. However, these operators are typically designed for continuous or discrete vectors and are not well-suited for handling natural language, such as recommendation prompts. Recent studies have demonstrated that, with appropriate prompts, LLMs can themselves serve as effective crossover and mutation operators for natural language processing tasks [12]. Following this approach, we employ an LLM-based crossover and mutation mechanism in this work. The prompt used for the LLM-based crossover and mutation operator is illustrated in Fig. 5. This prompt simulates the crossover and mutation process in conventional EAs. Specifically, the LLM performs crossover by combining two parent recommendation prompts, ensuring a diverse set of base recommendation prompts. Subsequently, LLM-based mutation employs conditional generation to fine-tune the newly generated prompts, further enhancing their diversity. Compared to directly fine-tuning the recommendation prompts, this approach tends to achieve better population diversity. Consequently, the LLM-based operator facilitates the effective generation of an offspring population from the parent population.
@@ IV-B1 LLM-NSGA-II
0: The population size NN; objective functions $\textbf{F}(\textbf{p})$;
0: The nondominated recommendation prompt set $\mathcal{P}$;
1: Generate the initial population $\mathcal{P}=\{\textbf{p}_{1},\ldots,\textbf{p}_{N}\}$ by using the initial operations introduced in Section IV-A1, and evaluate $\mathcal{P}$ by using the objective functions;
@@
3:while $t<t_{max}$ do
4:  Generate an offspring population of recommendation prompts $\mathcal{Q}=\{\textbf{q}_{1},\ldots,\textbf{q}_{N}\}$ according to the operations introduced in Section IV-A2;
5:  Utilizing the environmental selection of NSGA-II to select recommendation prompts from $\mathcal{P}\cup\mathcal{Q}$, thus forming the new population $\mathcal{P}$;
6:  $t\leftarrow t+1$;
7:end while
8:return A population of nondominated recommendation prompts $\mathcal{P}$.
@@ IV-B2 LLM-MOEA/D
0: The population size NN; objective functions $\textbf{F}(\textbf{p})$; set of decomposition weight vectors $\mathcal{W}=\{\textbf{w}_{1},\ldots,\textbf{w}_{N}\}$;
0: The nondominated recommendation prompt set $\mathcal{P}$;
1: Generate the initial population $\mathcal{P}=\{\textbf{p}_{1},\ldots,\textbf{p}_{N}\}$ by using the initial operations introduced in Section IV-A1, and evaluate $\mathcal{P}$ by using the objective functions;
2: Generate the neighborhood set $\mathcal{B}_{i}$, the external population $\mathcal{E}$, and the reference point $\textbf{z}=(z_{1},...,z_{m})^{T}$ following the MOEA/D framework;
@@
4:while $t<t_{max}$ do
5:  for $i=\{1,\ldots,N\}$ do
6:   Generate an offspring $\textbf{p}^{new}$ following the MOEA/D framework by using the LLM-based offspring generation introduced in Section IV-A2, and evaluate $\textbf{p}^{new}$ by using the objective functions;
7:   Update the reference point z, the population $\mathcal{P}$ and the external population $\mathcal{E}$ following the MOEA/D framework;
8:  end for
9:  $t\leftarrow t+1$;
10:end while
11:return An external population of nondominated recommendation prompts $\mathcal{E}$.
Decomposition-based evolutionary multiobjective optimization is a classical paradigm in multiobjective optimization. LLM-MOEA/D follows the decomposition-based method, with the pseudo-code presented in Algorithm 2. As in the MOEA/D framework, a neighborhood set $\mathcal{B}_{i}$ is initialized for each weight vector $\textbf{w}_{i}$. The initial population $\mathcal{P}$ is generated using the LLM-based initialization described in Section IV-A1, and all recommendation prompts in $\mathcal{P}$ are evaluated using the objective functions. Based on these evaluations, the external population $\mathcal{E}$ and the reference point z are initialized. During the iteration process, for each weight vector, an offspring recommendation prompt $\textbf{p}^{new}$ is produced using the LLM-based offspring generation method detailed in Section IV-A2. After evaluating $\textbf{p}^{new}$, the reference point z, the population $\mathcal{P}$, and the external population $\mathcal{E}$ are updated in accordance with the MOEA/D framework [39]. Once the termination conditions are satisfied, the external population $\mathcal{E}$ of nondominated recommendation prompts is returned.